# Learning Logical Rules using Minimum Message Length

### v1

@@ Abstract
Unifying probabilistic and logical learning is a key challenge in AI. We introduce a Bayesian inductive logic programming approach that learns minimum message length programs from noisy data. Our approach balances hypothesis complexity and data fit through priors, which explicitly favour more general programs, and a likelihood that favours accurate programs. Our experiments on several domains, including game playing and drug design, show that our method significantly outperforms previous methods, notably those that learn minimum description length programs. Our results also show that our approach is data-efficient and insensitive to example balance, including the ability to learn from exclusively positive examples.
Introduction
@@
Inductive logic programming (ILP) is a form of machine learning. The goal is to induce a logic program (a set of logical rules) that generalises examples and background knowledge (BK) (Muggleton 1991; Cropper and Dumancic 2022).
@@
As with all forms of machine learning, ILP learns from noisy data. A popular approach is to search for a minimal description length (MDL) (Rissanen 1978, 1998, 1987; Grünwald 2007) hypothesis (Quinlan 1990; Muggleton 1995; Srinivasan 2001; Cabalar, Pearce, and Valverde 2007; Hocquette et al. 2024). MDL is a mathematical formalisation of Occam’s razor: a preference for simpler (formalised as shorter) explanations of data. The goal of MDL is to encode data as a binary string as concisely as possible.
@@
Minimum message length (MML) (Wallace 2005) is another formalisation of Occam’s razor. The goal of MML is to find the most concise explanation of data, where an explanation is a two-part message consisting of a hypothesis (the first part) followed by the data given that the hypothesis is true (the second part).
@@
MDL and MML are closely related, but differ in philosophy, mathematical grounding, and practical usage (Baxter and Oliver 1994; Makalic and Schmidt 2022). We briefly summarise the key differences, but expand on them in Section 2. MML is explicitly Bayesian and uses priors over hypotheses and integrates them into encoding length. MDL avoids priors and often treats hypotheses as tools for encoding data rather than probabilistic entities. In MML, the message is two-part, first asserting the hypothesis, followed by data given that the asserted hypothesis is true. In MDL, the message may be one-part and often does not encode the hypothesis itself. MML and MDL also handle continuous parameters differently (Baxter and Oliver 1994; Grünwald 2007).
@@
In this paper, we use the MML principle to learn logic programs from noisy examples. Specifically, we design an encoding scheme which allows for lossless communication of a probabilistic logic program (De Raedt, Kimmig, and Toivonen 2007) and examples. The cost of a hypothesis is the length of the hypothetical message which would result if we were to communicate the examples using this hypothesis.
Our main claim is that MML outperforms a previous MDL-based score (Hocquette et al. 2024), shown to achieve state-of-the-art performance when learning logic programs from noisy examples. We term this previous score C-MDL. We argue that MML is better suited for this task for two reasons: (i) it uses prior probabilities, allowing us to favour more general programs, and (ii) it handles continuous parameters in a way that provides automatic regularisation (Wallace 2005; Farr and Wallace 2002; Wallace and Freeman 1987). MML has two more advantages over C-MDL: (iii) its hypothesis cost corresponds to an exact message length, and (iv) it uses a structured likelihood function that penalises over-generalisation, with this penalty being regulated by the generality priors from (i). Whilst an exact MDL approach could incorporate (iii) and (iv), it would still lack the intrinsic MML benefits of (i) and (ii).
@@
MML performs similarly to C-MDL on general tasks
@@
MML outperforms C-MDL when learning from different example proportions, including learning from exclusively one type of example
@@
MML requires less data than C-MDL
@@
    We formulate the program learning problem in Bayesian and MML terms. We use the MML framework to derive a cost function for learning probabilistic logic programs, consistent with the Problog semantics (De Raedt, Kimmig, and Toivonen 2007).
@@
ILP. Most ILP approaches learn from noisy data (Quinlan 1990; Muggleton 1995; Blockeel and De Raedt 1998; Srinivasan 2001). For instance, MaxSynth (Hocquette et al. 2024) searches for programs that minimise the total of the training errors and literals in the program. In contrast to our approach, these approaches do not allow for the explicit declaration of prior beliefs regarding the quantity of noise in the examples, do not employ techniques to mitigate the effect of example balance in their learning, and do not return probabilistic logic programs.
@@
Positive only learning. While several ILP approaches have been proposed that learn from exclusively positive examples (De Raedt and Bruynooghe 1993; Quinlan and Cameron-Jones 1995; Mooney and Califf 1995), most require an assumption of example completeness. One exception (Muggleton 1996) also models the program learning task under a Bayesian framework, but is restricted to noiseless data, uses coarse size-related priors, and employs a greedy search to find the hypothesis of maximum posterior probability. We differ by handling noise and using a prior that exactly corresponds to the program message length.
@@
MML and MDL. MML and MDL have been applied to learn a wide range of models, including neural networks (Makalic, Allison, and Dowe 2003; Hinton and Van Camp 1993), decision trees (Wallace and Patrick 1993), automata (Wallace 2005), and statistical distributions (Farr and Wallace 2002). The most relevant prior work is (Conklin and Witten 1994), which learned deterministic logic programs. However, their coding scheme is inefficient and cannot handle noisy data. We introduce a more principled coding scheme, an improved search procedure, and extend the framework to learn probabilistic logic programs. We choose MML over MDL for its Bayesian framework and method of handling parameter uncertainty: encoding real-valued parameters to a finite precision. MDL usually uses approximations or penalised maximum likelihood (ML) estimators to estimate continuous parameters. MML’s handling often yields less biased and consistent estimators that are often more accurate than ML (Baxter and Oliver 1994; Wallace 2005).
@@
Probabilistic logic programming. Previous work has focused on learning rule probabilities (Gutmann, Thon, and De Raedt 2011) or learning rules from Bayesian networks (De Raedt and Kersting 2008). Probfoil (De Raedt et al. 2015) is built on the Problog semantics (De Raedt, Kimmig, and Toivonen 2007) and learns probabilistic logic programs, but is frequentist, employs a greedy search, and requires both positive and negative examples. Markov logic networks (MLN) (Richardson and Domingos 2006) are another approach to unify probability and logic. MLNs are first-order knowledge bases with a weight attached to each logical formula and consider the joint distribution of a set of predicates. Our approach is Bayesian, uses a constraint solver in the search, can learn from exclusively positive examples, learns a single hypothesis, and employs MML as the hypothesis selection criterion.
Problem Setting
Inductive Logic Programming
We formulate our approach in the ILP learning from entailment setting (De Raedt 2008). We define an ILP input:
@@
We restrict the hypothesis space and background knowledge to definite programs with the least Hebrand model semantics (Lloyd 2012). We define a cost function:
@@
Given an ILP input $(E,B,\mathcal{H})$, a cost function $cost_{E,B}~:$$\mathcal{H}~\mapsto$$\mathbb{N}$ assigns a numerical cost to each hypothesis in $\mathcal{H}$.
Given an ILP input and a cost function $cost_{E,B}$, we define an optimal hypothesis:
@@
Given an ILP input $(E,B,\mathcal{H})$ and a cost function costE,B, a hypothesis $H\in\mathcal{H}$ is optimal with respect to costE,B when $\forall h^{\prime}\in\mathcal{H}$, costE,B($H$) $\leq$ costE,B($H^{\prime}$).
We describe the C-MDL cost function used by Hocquette et al. (2024):
@@
Given a hypothesis $H$ containing $size(H)$ literals which results in $fp(H)$ false positives and $fn(H)$ false negatives on training examples, the C-MDL cost of this hypothesis is: $cost_{C\text{-}MDL}(H)=size(H)+fp(H)+fn(H)$.
Next, we introduce an MML-based cost function.
Minimum Message Length
MML aims to find the most concise two-part explanation of examples $E$. An explanation consists of a hypothesis $H$ (first part) and examples given the hypothesis is true $E|H$ (second part). We briefly describe this communication format, following Wallace and Patrick (1993).
A hypothesis $H$ can optimally encode examples $E$ as a binary string of length:
$C(E|H)=-\log_{2}{P(E|H)}$ (1)
Where $P(E|H)$ is the probability of $E$ given that $H$ is true. This string is decodable to a receiver with knowledge of $H$. If $H$ is unknown, it must be communicated to the receiver. One way is to prepend the example string with a second binary string that communicates $H$ in a code agreed a priori. The length of this string is $C(H)=-\log_{2}{P(H)}$. The total message length is:

### v2

@@ Abstract
Unifying probabilistic and logical learning is a key challenge in AI. We introduce a Bayesian inductive logic programming approach that learns minimum message length hypotheses from noisy data. Our approach balances hypothesis complexity and data fit through priors, which favour more general programs, and a likelihood, which favours accurate programs. Our experiments on several domains, including game playing and drug design, show that our method significantly outperforms previous methods, notably those that learn minimum description length programs. Our results also show that our approach is data-efficient and insensitive to example balance, including the ability to learn from exclusively positive examples.
1 Introduction
@@
Inductive logic programming (ILP) is a form of machine learning. The goal is to induce a hypothesis (a set of logical rules) that generalises examples and background knowledge (BK) Muggleton (1991); Cropper and Dumancic (2022).
@@
As with all forms of machine learning, ILP handles noisy data. A popular approach is to search for a minimal description length (MDL) hypothesis Cabalar et al. (2007); Hocquette et al. (2024). MDL is a formalisation of Occam’s razor Rissanen (1978, 1998, 1987). The goal is to encode data as a binary string as concisely as possible.
@@
Minimum message length (MML) Wallace (2005) is another formalisation of Occam’s razor. The goal is to find the most concise explanation of data, where an explanation is a two-part message consisting of a hypothesis (the first part) followed by the data given that the hypothesis is true (the second part).
@@
MDL and MML are closely related, but differ in philosophy, mathematical grounding, and practical usage Baxter and Oliver (1994); Makalic and Schmidt (2022). We expand on the differences in Section 2, but briefly summarise them now. MML is explicitly Bayesian and uses priors over hypotheses and integrates them into the message length. MDL avoids priors and often treats hypotheses as tools for encoding data rather than probabilistic entities. In MML, the message is two-part, first asserting the hypothesis, then the data given that the hypothesis is true. In MDL, the message may be one-part and often does not encode the hypothesis itself. MML and MDL also handle continuous parameters differently Baxter and Oliver (1994); Grünwald (2007).
@@
In this paper, we use the MML principle to learn hypotheses from noisy examples. Specifically, we design an MML cost function for learning hypotheses expressed as probabilistic logic programs De Raedt et al. (2007). The cost of a hypothesis is the length of the message if we were to communicate the examples using this hypothesis.
Our main claim is that our MML method outperforms an existing MDL method Hocquette et al. (2024), shown to achieve state-of-the-art performance when learning logic programs from noisy examples. We call this previous score C-MDL.
Our MML method has three major advantages over C-MDL. First, the cost is the exact message length, resulting in more precise comparison between hypotheses. The increased precision enables us to determine when there is sufficient data to learn a more complex hypothesis, allowing us to generalise more reliably from sparse data. Second, our method uses a structured likelihood function that penalises over-generalisation, which enables learning from unbalanced, or one-class data. Third, MML uses prior probabilities which allows us to favour more general hypotheses. Favouring more general hypotheses prevents overfitting, as we do not wish to learn overly specific conclusions from small amounts of data
@@
MML performs similarly to C-MDL on noisy data and general tasks
@@
MML outperforms C-MDL when learning from different example proportions, including exclusively positive or negative examples
@@
MML is more data efficient than C-MDL
@@
    We formulate the ILP problem in Bayesian and MML terms. We use the MML framework to derive a cost function for learning probabilistic logic programs, consistent with the Problog semantics De Raedt et al. (2007).
@@ 2 Related Work
ILP. Most ILP approaches learn from noisy data Quinlan (1990); Muggleton (1995); Blockeel and De Raedt (1998); Srinivasan (2001). For instance, MaxSynth Hocquette et al. (2024) searches for programs that minimise the total of the training errors and literals in the program. In contrast to our approach, these approaches do not allow for the explicit declaration of prior beliefs regarding the quantity of noise in the examples, do not employ techniques to mitigate the effect of example balance in their learning, and do not return probabilistic logic programs.
@@
Positive only learning. Several ILP approaches learn from exclusively positive examples De Raedt and Bruynooghe (1993); Quinlan and Cameron-Jones (1995); Mooney and Califf (1995). Most assume example completeness. One exception Muggleton (1996) also learns in a Bayesian framework, but is restricted to noiseless data, assigns groups of hypotheses the same prior probability based only on size, and employs a greedy search to find the hypothesis of maximum posterior probability. We differ by handling noise, using a more sensitive prior tailored to each program, and utilising a SAT solver in our search procedure.
@@
MML and MDL. MML and MDL have been applied to learn a wide range of models, including neural networks Makalic et al. (2003); Hinton and Van Camp (1993), decision trees Wallace and Patrick (1993), automata Wallace (2005), and statistical distributions Farr and Wallace (2002). Conklin and Witten Conklin and Witten (1994) and Dowe et al. Dowe et al. (2007) both use MML to learn deterministic logic programs. Castillo-Andreu Castillo-Andreu (2012) uses MML to learn probabilistic programs. None of the cost functions in these works consider negative examples and do not use an MML-based procedure for estimation of continuous parameters. By contrast, our cost function considers negative examples and we use an MML-based procedure for estimation of continuous parameters. Moreover, we use predicate priors to express an explicit preference for more general programs and we use a constraint-based search procedure. We choose MML over MDL for its Bayesian framework and method of handling parameter uncertainty: encoding real-valued parameters to a finite precision. MDL usually uses approximations or penalised maximum likelihood (ML) estimators to estimate continuous parameters. MML’s handling often yields less biased and consistent estimators that are often more accurate than ML Baxter and Oliver (1994); Wallace (2005).
@@
Probabilistic logic programming. Previous work has focused on learning rule probabilities Gutmann et al. (2011) or learning rules from Bayesian networks De Raedt and Kersting (2008). Probfoil De Raedt et al. (2015) is built on the Problog semantics De Raedt et al. (2007) and learns probabilistic logic programs, but is frequentist, employs a greedy search, and requires both positive and negative examples. Markov logic networks Richardson and Domingos (2006) are first-order knowledge bases with a weight attached to each logical formula and consider the joint distribution of a set of predicates. Our approach is Bayesian, uses a constraint solver in the search, can learn from exclusively positive examples, learns a single hypothesis, and employs MML as the hypothesis selection criterion.
3 Problem Setting
3.1 Inductive Logic Programming
We define an ILP input in the learning from entailment setting De Raedt (2008):
@@ 3.1 Inductive Logic Programming
Background knowledge is a definite logic program. The hypothesis space contains Problog De Raedt et al. (2007) programs of a particular syntactic form:
Definition 2 (Probabilistic hypothesis).
Let $g$ be a Datalog program, $f$ be the predicate symbol in the head of the rules in $g$, $a$ be the arity of $f$, $\theta_{+}$ and $\theta_{-}$ be probabilities >> 0.5, and $g_{\phi}$ be $g$ where every instance of the predicate symbol $f$ is replaced with the predicate symbol $\phi$, where $\phi$ is a new predicate symbol of arity $a$. A probabilistic hypothesis $h=g_{\phi}\cup\{r_{+},r_{-}\}$, where $r_{+}$ and $r_{-}$ are:
$r_{+}$ $=\ \ \ \theta_{+}::$ $f(X_{1},\dots,X_{a})\leftarrow\phi(X_{1},\dots,X_{a})$
$r_{-}$ $=1-\theta_{-}::$ $f(X_{1},\dots,X_{a})\leftarrow\text{not }\phi(X_{1},\dots,X_{a})$
The function $\text{det}(h)=g$ maps a probabilistic hypothesis $h$ to its deterministic counterpart $g$. Informally, $h$ is a probabilistic wrapper around $\text{det}(h)$, and $h$ makes identical predictions to $\text{det}(h)$ but assigns probabilities to the predictions, governed by $\theta_{+}$ and $\theta_{-}$.
We define a cost function:
@@
Given an ILP input $(E,B,\mathcal{H})$, a cost function $\mathcal{C}_{E,B}~:~\mathcal{H}~\to~\mathbb{R}_{>0}$ assigns a numerical cost to each hypothesis in $\mathcal{H}$.
We define an optimal hypothesis:
@@
Given an ILP input $(E,B,\mathcal{H})$ and a cost function $\mathcal{C}_{E,B}$, an optimal hypothesis is $h^{*}=\arg\min_{h\in\mathcal{H}}\;\mathcal{C}_{E,B}(h)$.
Given a hypothesis $h$ and background knowledge $B$, a true positive is a positive example entailed by $det(h)\cup B$. A true negative is a negative example not entailed by $det(h)\cup B$. A false positive is a negative example entailed by $det(h)\cup B$. A false negative is a positive example not entailed by $det(h)\cup B$. For brevity, we omit $E$ and $B$ and denote the number of true positives, true negatives, false positives, and false negatives of hypothesis $h$ as $tp(h)$, $tn(h)$, $fp(h)$, and $fn(h)$ respectively.
We describe the C-MDL cost function used by Hocquette et al. Hocquette et al. (2024)
@@
Let $h$ be a hypothesis with $fp(h)$ false positives and $fn(h)$ false negatives on training examples, and $size(g)$ denote the number of literals in the definite program $g$. The C-MDL cost of $h$ is $\text{C-MDL}_{E,B}(h)=size(\text{det}(h))+fp(h)+fn(h)$.
We describe a notion of entailed atoms and entailed examples with respect to a hypothesis
Definition 6 (Entailed Examples).
Given a hypothesis $h$ and an instance space $X$ containing all possible examples, we define a set of entailed atoms as $\{x\in X\;|\;B\cup det(h)\models x\}$ and a set of non-entailed atoms as $\{x\in X\;|\;B\cup det(h)\not\models x\}$. Given a set of examples $E\subseteq X$, we define a set of entailed examples as $\{e\in E\;|\;B\cup det(h)\models e\}$ and a set of non-entailed examples as $\{e\in E\;|\;B\cup det(h)\not\models e\}$.
4 Minimum Message Length
Given a prior and a likelihood function, MML aims to find the most concise two-part explanation of examples $E$. An explanation consists of a hypothesis $h$ (first part) and the examples given that hypothesis is true $E|h$ (second part). We briefly summarise this communication format following Wallace and Patrick Wallace and Patrick (1993).
Given a hypothesis $h$, the examples $E$ can be optimally encoded as a binary message of length:
$C(E|h)=-\log_{2}P(E|h)$
where $P(E|h)$ is the probability of observing $E$ given that $h$ is true. This message can be decoded by a receiver who already knows $h$. If the receiver does not know $h$, the sender must also send $h$, using a code agreed upon a priori. The length of this message is:
$C(h)=-\log_{2}P(h)$
The total message length required to send the hypothesis and the examples is therefore: